Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: Running established version of M&C/data acquisition system; did have a substantial crash on Saturday morning which took many hours to recover from; still being investigated (CBE node went down; many files in the Executor queue complicated the recovery; new procedure to try out). In the final week of D configuration.
  • New proposal cycle advertised in e-News; OSRO/RSRO/OSS updated for call.
  • Thursday - tentative detailed discussion on 3-bit status (Perley+).
  • Key dates:
    • 13 Sep: Begin move to DnC configuration
    • 13-17 Sep: 74 MHz array testing campaign
    • 15 Sep: Call for Proposals
      • Note: Ku/X-wide will not be offered in main call (allowed for exploratory).
    • 01 Oct: Proposal deadline (A config; lowband for end of B config)
    • 04 Oct: Begin move to C configuration
    • 13-15 Oct: Phased Array testing campaign
  • Staff Travel:
    • Chandler: 01 Sep - 15 Sep
    • DRAO: Carlson: 16 Jul - 19 Aug
    • Rupen: 11 Sep - 21 Sep
    • Sowinski: 20 Sep - 15 Oct
    • McMullin: 01-10 Oct (vacation/Chile)
    • van Moorsel: 11-15 Oct
    • Wrobel: 24 Oct-27 Oct
    • DRAO: Dougherty: 11 Oct - 12 Oct (SKA; ABQ)
  • Visitors:
    • Brent Carlson: 18 October - 25 October; second trip post SKA meeting TBD
    • DRAO: Dougherty: 12 Oct - 17 Oct (NRAO-DSOC)
    • Adam Deller/Joe Lazio: 13-15 Oct (phased array testing)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Notes: Antennas 18 and 26 have had cable changes to the L302s - please confirm presence/absence of phase jumps.
      • Antennas with 3-bit samplers: EA15, EA22, EA27, EA28
      • Antennas with 4-band dipoles: EA03, EA11 (East), EA01, EA19 (West), EA06, EA28 (North)
    • Links:
  • WIDAR (Rupen):
    • Notes: Plans for full population of BlBs/switch for October 1; software (CBE/CM) have been stable; some issues encountered with 3-bit testing that are being explored.

3-bit testing status

  • Notes: Possible update on Thursday; hunting down issues with some of the deployed samplers; now looking like the hardware itself might be the source of the issues with the noise characteristics. We will not meet our timeline for the order based on the difficulties so there will be a delay in the full deployment available for RSRO. Stay tuned.

Software status and updates (Butler)

  • Links:
  • Notes: Enjoyed a relatively stable period with no development last week; the CM/CBE averaging (in time) has risen in priority and will be worked on in the coming week; hoping for deployment by C configuration if testing time is adequate.
  • CASA Update (new functionality available/issues/enhancements). Flagging slowdown(!) introduced in last casapy-test; fixed as of Monday (06Sep10). Development cycle raps up next week; release in mid-October.
  • AIPS Update (new functionality available/issues/enhancements). Added PDVER parameter back to many tasks (allows choosing D term spectral table); note this could confuse use of tgets with prior versions. Working with Bill on OBIT filler; handles simple SDM files; it was noted there was a problem with updates in Socorro - being looked at. Also looking at applying opacity corrections for the CL table. Note: OBIT update problem has been resolved as of 8Sep.

Observing recommendations update (Ops)

Commissioning priorities for the week (McMullin, Chandler)

  • 1) RSRO/ECSO science
  • 2) 3-bit sampler testing
  • 3) OSRO/RSRO blockers (focused on missing data/data quality issues)
  • 4) CSV testing
  • 5) Demo science observing
  • 6) OSRO science observing

Commissioning Reports

Polarization: Taylor: Polarization calibration test - delayed.

Spectral Line: Dhawan: Frequency offset tests (for Doppler tracking) - delayed.

Switched power update (McMullin): Meeting last week that hashed out the unified use of SDM -> MS -> UVFITS tables for calibration. EVLA-specific SysCal table is being written now; updated as of today with fixes from Steve's review. Proposed use of CalDevice table (being written out currently in dummy form). Hoped to have filler writing these bits by next week.

In preparation

  • Polarization: Multi-band leakage/PA (Myers, Marvil): Needed.
  • Spectral Line: Momjian: Zeeman calibration test.
  • Spectral Line: Bandpass stability,multi-band on 3C84 et al. (Goss/Chomiuk);

Commissioning Test Planning

  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Delayed
  • Other:

RSRO/ECSO Reports

Project Band(s) Test Code Test Status Allocation Status (D) Notes JIRA
RSRO            
AA30 Q TRSR0003 - 0/51 Ready
1773293 (5.5 hr; 0630-0830)
RSRO-1
AB1345 L TRSR0004 21May10 23/40 In Progress;Needs SB RSRO-2
AB1346 K TRSR0005
TSPE0001
23Jun10 24/42 Ready
1859763 (3.5 hr; 1630-1730)
RSRO-3
AC982 CKKaQ TRSR0002 28Jul10 30/30 Complete; pending data review RSRO-4
AH1006 LCXKa TRSR0006 - 16/36 In Progress; Need SBs RSRO-5
AK726 Ka TRSR0007 - 24/29.5 In Progress; Need SBs RSRO-6
AL746 CKa TRSR0008 05Jul10 11/11 Complete RSRO-7
AM1014 C TRSR0009 22Jun10 0/0 C config RSRO-9
AS1015 C TRSR0010 21May10 4/4 Complete RSRO-11
AT374 C TRSR0011 25Jul10 0/0 C config RSRO-12
AV318 L TRSR0012 - 12/12 Complete; PI converted to OSRO RSRO-13
ECSO          
AW775 L TRSR0017 05Jul10 56/72 In Progress
1660006 (7)
-
AM1005 KKaQ TRSR0014 - 3/20 In Progress
1833848 (3 hr; 1830-1830)
-
AM1038 Q TRSR0019 21May10 10/10 Complete -
AM1040 Ka TRSR0019 21May10 10/10 Complete -
AC992 L TRSR0018 26Jun10 4/4 Complete -
AO255 Ka TRSR0016 - 36/40 In progress
1756969 (4)
-
AM1018 Ka TRSR0015 - 0/48 No SBs submitted -
AB1353 Ka TRSR0013 - 0/51 No SBs submitted -
  • Notes:
    • Current Residents
      • Laura Perez (POC: Claire Chandler); area(s): High frequency observing
      • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).
      • Volker Heesen (POC: Michael Rupen); area(s): polarimetry
      • Peter Hofner (POC: TBD); area(s): post-processing
      • Russ Taylor (POC: Rick Perley); area(s): polarimetry
      • NewSatoko Takahashi (POC: Juergen Ott); area(s): continuum, mosaicing; 10B-126 (AT380): A Millimeter Study of the Embedded Star-Cluster in the Orion Molecular Cloud.
        • NewHauyu Liu (POC: Juergen Ott); area(s): continuum, mosaicing
      • NewGreg Hallinan (POC: TBD); area(s):; 10B-209 (AH1020): Broadband Periodic Dynamic Spectra of Ultracool Dwarf Pulsars
  • Notes: AB1346 - issues with performance/data sizes with CASA - being investigated; problems with specific spw in their setup (Michael has analyzed this and found that the problems are in line with expectations based on the behavior of both the baseband and subband edges. AM1038 - reported on timings for different steps (import+=12 hours; asdm2MS=6 hours; flagging=3 hours, etc); noted 3 missing scans (one mislabeled, two no data). Noted issues with specific antennas/spws but overall data okay. AH1006: further progress - getting down to predicted sensitivity (17microJy/beam); more SBs will be submitted. AK726: struggling with data volume (as are all RSRO); working through last data sets; will complete D config allocations.

Demo Science (Chandler/Condon): Project Codes: *TDEM*

    • Status: Begun; list requires prioritization; limited remaining time in D-config; Will try for IRC+10216 this week as possible.
    • Testing
Project Status Time Comments SBSorted ascending
Crab nebula Complete 10/10 Complete; L-band OSRO1/OSRO2  
IRC+10216 Ready (HF) 8/33 In progress; K, Ka, Q spectral sweep; no recirculation; pending test  
M51 Ready (MR) 1.5/8 In progress; L band; needs corr test  
M87 Ready (HF) 0/6 Ka TDEM0006_sb1470572 (not submitted; 6 hr: 0900-1000 LST)
Herc A Ready 3/5 CX bands TDEM0011_sb1729424 (5 hr; 1500-1700)
M82 Complete 2/2 L band Complete
Cygnus A Ready 3x6 1x10 possible; 12 OSRO hours obtained Withdrawn

IRC+10216 Update report (Claussen)

  • No exciting spectra yet; into the OSRO focus phase for observations. More later.

Commissioning Group updates

User Support Update (van Moorsel) - push to next week.

AOB

Misc Actions

  • [McMullin]: Splatalogue use for RFI line list.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin/Butler/Rupen/Owen]: Resolve discussion of requirements for DnC test of 4-band.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 02 Sep 2010
Topic revision: r6 - 2010-09-10, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback